The Raymond Burr National Bank has $1,000 in assets with an average duration of 5 years. This bank has $800 in liabilities with an average duration of 6.25 years. Market interest rates start at 6 percent and fall by 1 percent. What is the change in net worth of this bank?

Please show the steps to solve this without using Excel.
Answer #1

increase in asset = -asset duration*change in yield*asset value

=-5*(-0.01)*1000=50

increase in liabilities = =6.25*(-0.01)*800=50

change in networth = increase in asset-increase in liabilities = 50-50 = 0
Answer #2

Calculating the Change in Net Worth of the Bank

Given:

  • Assets:

    • Value = $1,000

    • Duration (DA) = 5 years

  • Liabilities:

    • Value = $800

    • Duration (DL) = 6.25 years

  • Interest Rate Change (Δr): Decrease of 1% (from 6% to 5%).


Step 1: Understand the Duration Gap Formula

The change in net worth (ΔNW) due to interest rate changes is calculated using:

ΔNW(DADL×LA)×A×Δr1+r

Where:

  • A = Assets ($1,000)

  • L = Liabilities ($800)

  • r = Initial interest rate (6% or 0.06)

  • Δr = Interest rate change (-1% or -0.01).


Step 2: Plug in the Values

  1. Calculate the Duration Gap:

    Dgap=DADL×LA=56.25×8001000=55=0

    (The duration gap is zero, meaning the bank is initially immunized against interest rate changes.)

  2. Compute Change in Net Worth:

    ΔNW0×1000×0.011.06=0


Answer:

The change in net worth is $0.


answered by: anonymous